OpenStudy (anonymous):

Giving medals solve 2x-5/x+4x-1/x+2=-3x+8/x^2+2x a. -1+-radical433/12 b. 2, -3/2 c. -2/9 d. -2/3, 1/2

OpenStudy (anonymous):

do u mean this? 2x-(5/x)+4x-(1/x)+2=-3x+(8/x^2)+2x

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

yikes, haha

OpenStudy (anonymous):

do you know how to do it?!

OpenStudy (anonymous):

I can help, yes. first thing is we want to clear the fractions, look for a polynomial we can multiply both sides by that can clear the fractions

OpenStudy (anonymous):

x+2

OpenStudy (anonymous):

that was my first pick aswell ^_^ that would clear the fractions on the left hand side, but not the fraction on the right hand side. so, we have to use (x^2+2x) and multiply it to both sides

OpenStudy (anonymous):

can you show me?! lol

OpenStudy (anonymous):

ya, 1 sec then...

OpenStudy (anonymous):

so on the right we have: \[\frac{ 3x+8 }{ x^2+2x } * (x^2+2x) \rightarrow 3x+8\] '[\frac{ 3x+8 }{ x^2+2x } * (x^2+2x) \rightarrow 3x+8\]

OpenStudy (anonymous):

the first term on the left would look like this: \[\frac{ 2x-5 }{ x }*(x^2+2x)\rightarrow (2x-5)(x+2)\] '[\frac{ 2x-5 }{ x }*(x^2+2x)\rightarrow (2x-5)(x+2)\]

OpenStudy (anonymous):

wait so is it C?

OpenStudy (anonymous):

do you wanna try the second term on the left? \[\frac{ 4x-1 }{ x+2 }*(x^2+2x)\rightarrow ?\]

OpenStudy (anonymous):

are you over working this problem out? it is kinda long if i'm going to show you each step...

OpenStudy (anonymous):

sorta. what do you think it is?! lol

OpenStudy (anonymous):

I get that way too... what do I think it is.... I think it will have 2 solutions. have you plugged in any of the solutions to see if the statment is true?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

haha, idk! we can work it out some more together if u want tho... ur choice

OpenStudy (anonymous):

PLEASE!! lol i need help

OpenStudy (anonymous):

did you not solve it all the way yet?!

OpenStudy (anonymous):

no, not yet

OpenStudy (anonymous):

here, you do this part \[\frac {4x−1}{x+2}∗(x^2+2x)\rightarrow?\] '[\frac {4x−1}{x+2}∗(x^2+2x)\rightarrow?\]

OpenStudy (anonymous):

or is it easier if I write it this way: \[(4x−1)* \frac {x^2+2x}{x+2}\rightarrow?\]

OpenStudy (anonymous):

wait i got 4x-1.... lol

OpenStudy (anonymous):